# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v2

@@ Appendix B Implementation details
The Prob-PSENN autoencoder module has been implemented as a CNN-based Variational Autoencoder, consisting of four convolutional layers with $4\times 4$ kernels, batch normalization, and LeakyReLU activations, followed by a $1\times 1$ convolutional layer that controls the dimensionality of the latent space. The decoder mirrors this architecture by means of deconvolutional layers. Further details on the architecture and the corresponding hyperparameters are provided in Figure B.12. The inputs values have been normalized to the range $[-1,1]$, and, for grayscale datasets, the input shapes have been upscaled from $28\times 28$ to $32\times 32$ using bilinear interpolation. For the SVHN dataset, random brightness adjustments and zoom transformations have been incorporated during training to improve generalization. The experiments have been run for $\tau_{1}=1.0,\tau_{2}=10.0,\tau_{3}=0.05,\tau_{4}=0.05$ and a batch size of 128. The models have been trained for 100 epochs on grayscale datasets and 30 epochs on the SVHN dataset. The Adam optimization method [22] has been used to train the networks, using a learning rate of $0.001$. The final classification layer $W$ has been set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal. In order to ensure higher stability during the training process, the value of $\mathscr{L}_{INT}$ has been cut to the range $[-1,10^{3}]$. For all datasets, standard training and test partitions have been used, with 10% of the training data reserved for validation. The results of [27, 13] have been reproduced using the implementations released by the authors. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
@@
Figure B.12: VAE architectures used in Prob-PSENNs for (a) MNIST, F-MNIST, K-MNIST, and (b) SVHN.
@@ Appendix C Complementary results on the MNIST dataset
In order to complement the results reported in Section 4, Figure C.13 shows, for the MNIST dataset, 15 randomly sampled sets of prototypes from a Prob-PSENN with $l=5$ (left) and $l=10$ (right). As can be observed, the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, as well as to obtain more detailed decodings.
@@ Appendix D Results on the Fashion-MNIST dataset
Furthermore, a visual comparison between the explanations provided by a non-probabilistic PSENN and a Prob-PSENN is provided in Figures D.15 and D.16. For the sake of completeness, the former figure includes the explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) prototypes. In both cases, the original input is shown in the top-left corner, and the prototypes are colored according to the classes returned by the model when the prototypes are classified, in order to make it clearer which class each prototype is most associated with. Notice that this connection is explicit and unequivocal by construction in our Prob-PSENNs, being the class $y_{i}$ represented by the prototype $r_{i}$, $1\leq i\leq c$, resulting in more transparent decision processes and easier-to-interpret explanations. As can be observed in Figures D.15 and D.16, the prototypes learned by the PSENNs are, even for $m=50$ prototypes, considerably less meaningful and varied than the ones captured by Prob-PSENN. Furthermore, the uncertainty metrics introduced in Section 3.3 enable us to quantify a low aleatoric and epistemic explanatory uncertainty for this input ($0.004$ and $0.16$, respectively), which increases the trustworthiness and reliability of the outcomes of Prob-PSENN.
@@ Appendix E Results on the K-MNIST dataset
This section contains additional results on the 10-class K-MNIST dataset [8], which includes handwritten representations of Japanese letters. Random images from this dataset are shown in Figure E.17, from which it can be noticed a high variance in the representation of each letter. Randomly sampled prototypes from Prob-PSENNs with $l=5$ and $l=10$ are displayed in Figure E.18. As can be assessed in the figure, despite the complexity and high variability of this dataset, Prob-PSENNs are capable of capturing the most frequent representations of each class, ensuring a sufficiently diverse and meaningful characterization of the classes.
As for the previous datasets, a comparison between the explanations of PSENNs and Prob-PSENN for the K-MNIST dataset is provided in Figures E.19 and E.20. Notice how, also due to the higher variability in the representations of each class, the prototypes learned by a PSENN with $m=15$ prototypes (Figure E.19-left) are not sufficiently representative of the classes, and, consequently, we need to resort to a large number of prototypes ($m=50$, Figure E.19-right). However, even for $m=50$, the prototypes learned by PSENN are considerably less varied and informative than the ones sampled by Prob-PSENN, which evidences the higher flexibility of our approach, as well as its increased transparency and explanatory capabilities.
@@ Appendix F Ablation studies: complementary results
This Appendix complements the results from Section 4.4 by comparing the performance of Prob-PSENN with its deterministic counterpart, using the same architecture and training pipeline. The point-wise prototypes have been optimized using the loss function proposed in [27]. The results obtained with the deterministic PSENN are summarized in Table F.1, computed considering $l=20$ and $l=40$ dimensions for the latent space and various numbers of prototypes $m$. The results for Prob-PSENN are included in the last rows of the table (for $l=20$ and $l=40$), as reference. As can be observed, the performance of both approaches is nearly identical, with differences in average accuracy below 0.2%. These results further demonstrate that the advantages of our probabilistic model do not imply a reduction in its predictive performance, while bringing as benefits enhanced trustworthiness, uncertainty quantification and explanatory capabilities.
@@
Table F.1: Predictive accuracy percentages (mean and standard deviation for 5 runs) for deterministic PSENNs using the same architecture and training pipeline as Prob-PSENNs. The best results for each approach (PSENN and Prob-PSENN) are highlighted in bold, in order to facilitate comparison and emphasize their similar performance.
Model Configuration MNIST F-MNIST K-MNIST
PSENN ($m\!=\!10$, $l\!=\!20$) 99.39 $\pm$ 0.04 92.11 $\pm$ 0.10 96.71 $\pm$ 0.12
PSENN ($m\!=\!10$, $l\!=\!40$) 99.42 $\pm$ 0.02 91.94 $\pm$ 0.25 96.79 $\pm$ 0.08
PSENN ($m\!=\!30$, $l\!=\!20$) 99.33 $\pm$ 0.10 90.92 $\pm$ 0.14 96.42 $\pm$ 0.18
PSENN ($m\!=\!30$, $l\!=\!40$) 99.29 $\pm$ 0.08 90.30 $\pm$ 0.51 96.72 $\pm$ 0.11
Prob-PSENN ($m\!=\!10$, $l\!=\!20$) 99.42 $\pm$ 0.05 91.91 $\pm$ 0.14 96.72 $\pm$ 0.21
Prob-PSENN ($m\!=\!10$, $l\!=\!40$) 99.38 $\pm$ 0.02 91.98 $\pm$ 0.24 96.61 $\pm$ 0.17
@@
Next, in Table F.2 we report the results obtained with Prob-PSENNs for which VAEs have been replaced with vanilla AutoEncoders, matching all the remaining architectural details and the training pipeline. The results show that vanilla AutoEncoders lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.21 visualizes different sets of randomly sampled prototypes from Prob-PSENNs with different backbone architectures, highlighting the advantage of regularized latent spaces for improving prototype quality in high-dimensional settings.
@@ Appendix G Results on the E-MNIST dataset
In order to analyze the scalability of Prob-PSENN in tasks consisting of larger number of classes, we evaluated its performance in the E-MNIST dataset [9] (47 classes), an extension of MNIST which includes both numeric and alphabetic characters. As it can be seen in Figure G.22, Prob-PSENN is still capable of learning suitable distribution over the prototypes for the 47 classes, yielding realistic and diverse prototypes, while achieving a test accuracy of 0.89.

### v3

@@ Appendix B Implementation details
For image datasets, the Prob-PSENN autoencoder module has been implemented as a CNN-based Variational Autoencoder, consisting of four convolutional layers with $4\times 4$ kernels, batch normalization, and LeakyReLU activations, followed by a $1\times 1$ convolutional layer that controls the dimensionality of the latent space. The decoder mirrors this architecture by means of deconvolutional layers. For tabular data, the architectures consist of three-layer MLPs with batch normalization and LeakyReLU activations. Further details on the architecture and the corresponding hyperparameters are provided in Figure B.13. The pixel values have been normalized to the range $[-1,1]$, and, for grayscale datasets, the input shapes have been upscaled from $28\times 28$ to $32\times 32$ using bilinear interpolation. For the SVHN dataset, random brightness adjustments and zoom transformations have been incorporated during training to improve generalization. The experiments have been run for $\tau_{1}=1.0,\tau_{2}=10.0,\tau_{3}=0.05,\tau_{4}=0.05$ and a batch size of 128. The models have been trained for 100 epochs on grayscale datasets and 30 epochs on the SVHN dataset. The Adam optimization method [23] has been used to train the networks, using a learning rate of $0.001$. The final classification layer $W$ has been set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal. In order to ensure higher stability during the training process, the value of $\mathscr{L}_{INT}$ has been cut to the range $[-1,10^{3}]$. For all datasets, standard training and test partitions have been used, with 10% of the training data reserved for validation. The results of [28, 14] have been reproduced using the implementations released by the authors. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
@@
(c) Sensorless Drive Diagnosis
Figure B.13: VAE architectures used in Prob-PSENNs for (a) MNIST, F-MNIST, K-MNIST, (b) SVHN and (c) Sensorless Drive Diagnosis.
@@ Appendix C Complementary results on the MNIST dataset
In order to complement the results reported in Section 4, Figure C.14 shows, for the MNIST dataset, 15 randomly sampled sets of prototypes from a Prob-PSENN with $l=5$ (left) and $l=40$ (right). As can be observed, the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, as well as to obtain more detailed decodings.
@@ Appendix D Results on the Fashion-MNIST dataset
Furthermore, a visual comparison between the explanations provided by a non-probabilistic PSENN and a Prob-PSENN is provided in Figures D.16 and D.17. For the sake of completeness, the former figure includes the explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) prototypes. In both cases, the original input is shown in the top-left corner, and the prototypes are colored according to the classes returned by the model when the prototypes are classified, in order to make it clearer which class each prototype is most associated with. Notice that this connection is explicit and unequivocal by construction in our Prob-PSENNs, being the class $y_{i}$ represented by the prototype $r_{i}$, $1\leq i\leq c$, resulting in more transparent decision processes and easier-to-interpret explanations. As can be observed in Figures D.16 and D.17, the prototypes learned by the PSENNs are, even for $m=50$ prototypes, considerably less meaningful and varied than the ones captured by Prob-PSENN. Furthermore, the uncertainty metrics introduced in Section 3.3 enable us to quantify a low aleatoric and epistemic explanatory uncertainty for this input ($0.004$ and $0.16$, respectively), which increases the trustworthiness and reliability of the outcomes of Prob-PSENN.
@@ Appendix E Results on the K-MNIST dataset
This section contains additional results on the 10-class K-MNIST dataset [9], which includes handwritten representations of Japanese letters. Random images from this dataset are shown in Figure E.18, from which it can be noticed a high variance in the representation of each letter. Randomly sampled prototypes from Prob-PSENNs with $l=5$ and $l=40$ are displayed in Figure E.19. As can be assessed in the figure, despite the complexity and high variability of this dataset, Prob-PSENNs are capable of capturing the most frequent representations of each class, ensuring a sufficiently diverse and meaningful characterization of the classes.
As for the previous datasets, a comparison between the explanations of PSENNs and Prob-PSENN for the K-MNIST dataset is provided in Figures E.20 and E.21. Notice how, also due to the higher variability in the representations of each class, the prototypes learned by a PSENN with $m=15$ prototypes (Figure E.20-left) are not sufficiently representative of the classes, and, consequently, we need to resort to a large number of prototypes ($m=50$, Figure E.20-right). However, even for $m=50$, the prototypes learned by PSENN are considerably less varied and informative than the ones sampled by Prob-PSENN, which evidences the higher flexibility of our approach, as well as its increased transparency and explanatory capabilities.
@@ Appendix F Ablation studies: complementary results
This section complements the results from Section 4.4 by comparing the performance of Prob-PSENN with its (i) black-box and (ii) deterministic counterparts. The black-box models use the same encoder architecture as Prob-PSENN, followed by a linear classifier, and have been trained end-to-end with the cross-entropy loss. For the sake of completeness, and to ensure a fair comparison, we also evaluated a black-box variant using the full VAE backbone (encoder and decoder), matching the regularization and training dynamics of Prob-PSENN’s backbone. Nonetheless, notice that the decoder plays no functional role in the black-box classification process at inference time. The deterministic PSENNs share the same architecture as Prob-PSENN, but replace the probabilistic prototype distributions with a fixed set of $m$ trainable point-wise prototypes. These models have been trained using the classification and reconstruction loss of Prob-PSENN, as well as the interpretability regularizers proposed in [28]. Both ablation variants have been evaluated considering $l=20$ and $l=40$ dimensions for the latent space, and various numbers of prototypes $m$ for the deterministic PSENNs.
@@
The results obtained with these ablated models are summarized in Table F.1. First, we can notice that Prob-PSENN matches or outperforms these black-box counterparts, a result consistent with prior related works [28, 14]. We attribute this to the structured clustering induced by the prototype distributions in the latent space and the robustness gained through averaging predictions over multiple sampled prototypes, reinforcing the practical advantages of Prob-PSENN’s pipeline. Next, comparing Prob-PSENN with their deterministic versions, it can be observed that the performance of both approaches is nearly identical, with differences in average accuracy below 0.2%. These results further demonstrate that the advantages of our probabilistic model do not imply a reduction in its predictive performance, while bringing as benefits enhanced trustworthiness, uncertainty quantification and explanatory capabilities.
@@
Table F.1: Predictive accuracy percentages (mean and standard deviation for 5 runs) for black-box and deterministic PSENNs using the same architecture and training pipeline as Prob-PSENNs. The best results for each approach are highlighted in bold, in order to facilitate comparison and emphasize their similar performance.
Model Configuration MNIST F-MNIST K-MNIST SVHN
Black-box (Enc. only) ($m=0$, $l\!=\!20$) 99.11 $\pm$ 0.19 90.77 $\pm$ 0.15 95.58 $\pm$ 0.57 90.94 $\pm$ 0.30
Black-box (Enc. only) ($m=0$, $l\!=\!40$) 99.00 $\pm$ 0.31 90.59 $\pm$ 0.22 95.63 $\pm$ 0.43 91.10 $\pm$ 0.39
Black-box (Enc. + Dec.) ($m=0$, $l\!=\!20$) 99.20 $\pm$ 0.03 91.66 $\pm$ 0.20 95.83 $\pm$ 0.16 91.53 $\pm$ 0.56
Black-box (Enc. + Dec.) ($m=0$, $l\!=\!40$) 99.17 $\pm$ 0.04 91.82 $\pm$ 0.16 95.93 $\pm$ 0.18 91.74 $\pm$ 0.41
PSENN ($m\!=\!10$, $l\!=\!20$) 99.39 $\pm$ 0.04 92.11 $\pm$ 0.10 96.71 $\pm$ 0.12 92.37 $\pm$ 0.28
PSENN ($m\!=\!10$, $l\!=\!40$) 99.42 $\pm$ 0.02 91.94 $\pm$ 0.25 96.79 $\pm$ 0.08 92.44 $\pm$ 0.32
PSENN ($m\!=\!30$, $l\!=\!20$) 99.33 $\pm$ 0.10 90.92 $\pm$ 0.14 96.42 $\pm$ 0.18 91.05 $\pm$ 0.41
PSENN ($m\!=\!30$, $l\!=\!40$) 99.29 $\pm$ 0.08 90.30 $\pm$ 0.51 96.72 $\pm$ 0.11 90.52 $\pm$ 0.23
Prob-PSENN ($m\!=\!10$, $l\!=\!20$) 99.42 $\pm$ 0.05 91.91 $\pm$ 0.14 96.72 $\pm$ 0.21 91.62 $\pm$ 0.56
Prob-PSENN ($m\!=\!10$, $l\!=\!40$) 99.38 $\pm$ 0.02 91.98 $\pm$ 0.24 96.61 $\pm$ 0.17 92.14 $\pm$ 0.24
@@
Finally, in Table F.2 we report the results obtained with Prob-PSENNs for which VAEs have been replaced with vanilla AutoEncoders, matching all the remaining architectural details and the training pipeline. The results show that vanilla AutoEncoders lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.22 visualizes different sets of randomly sampled prototypes from Prob-PSENNs with different backbone architectures, highlighting the advantage of regularized latent spaces for improving prototype quality in high-dimensional settings.
@@ Appendix G Results on the E-MNIST dataset
In order to analyze the scalability of Prob-PSENN in tasks consisting of a larger number of classes, we evaluated its performance in the E-MNIST dataset [10] (47 classes), an extension of MNIST that includes both numeric and alphabetic characters. As can be seen in Figure G.23, Prob-PSENN is still capable of learning suitable distributions over the prototypes for the 47 classes, yielding realistic and diverse prototypes, while achieving a test accuracy of 0.89.